OpenStudy (anonymous):

Is each situation a linear or a nonlinear model? Column A Column B 1. The stopping distance for a car is given by the equation , where D is the distance a car travels once the brakes are applied and v is the car’s velocity. 2. A home’s value is given by the equation y = 154,000(1.04)x, where y is the value of the house and x is the number of years since the home was purchased.

OpenStudy (anonymous):

can you help with 1 first i think its linear, right?

OpenStudy (solomonzelman):

are you sure you aren't missing any equations?

OpenStudy (anonymous):

3. The amount of money Elijah has in his bank account is given by the equation B = 1400 + 120m, where B is the balance in his account and m is the number of months since he opened the account. 4. The number of words Sarah can type in a given amount of time is given by the equation w = 32m, where w is the number of words and m is the number of minutes she types.

OpenStudy (anonymous):

there's the other ones sorry

OpenStudy (solomonzelman):

there is no equation for #1

OpenStudy (anonymous):

sorry

OpenStudy (solomonzelman):

okay, what can you tell me (about any of the questions you posted) so far?

OpenStudy (solomonzelman):

linear means a line, in a form of \(\large\color{black}{ y=mx+b }\) (where m and b are just numbers)

OpenStudy (solomonzelman):

~ when you have \(\large\color{black}{ y=x^c }\) and c is not 1 or 0, then it is not linear. ~ when you have \(\large\color{black}{ y=c^x }\) that is an exponential (unless x is 1 or 0)

OpenStudy (solomonzelman):

number 1, you have: \(\large\color{black}{ D=\frac{\LARGE 1}{\LARGE15.7 }v^2 }\) which is same as: \(\large\color{black}{ y=\frac{\LARGE 1}{\LARGE15.7 }x^2 }\)

OpenStudy (solomonzelman):

is number 1 a linear function?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@sleepyjess

OpenStudy (vickyisthesmartone123):

@TheSmartOne

TheSmartOne (thesmartone):

I got tagged to this 10 months ago, and now again? ;-;

OpenStudy (vickyisthesmartone123):

i need urt help?